An investor purchases 200 shares of a company at a purchase price of $30.00 at the start of the year. During the year, the company paid out $1.75 of dividends per share. The investor then sells all the shares at a selling price of $27.00. Determine the investor's total percentage return. 


Question 29 options:

-4.17%


-5.55%


-6.89%


-7.38%


None of the above
Answer #1

А 1 Purchase cost 2 Dividend paid 3 Selling 4 5 Total percentage return B $6,000.00 $ 350.00 $5,400.00 -4.17%

Calc:

A B 1 Purchase cost =200*30 2 Dividend paid =1.75*200 3 Selling =200*27 4 5 Total percentage return =(B3+B2-B1)/B1